OpenStudy (anonymous):

(√3-√5)(√3+√5) what is the product solve and show work

OpenStudy (anonymous):

\((a-b)(a+b)=a^2-b^2\)

OpenStudy (anonymous):

in your case \(a=\sqrt{3},a^2=3,b=\sqrt{5},b^2=5\) so you get \(3-5\)

OpenStudy (anonymous):

@satellite73 what the next step